OpenStudy (anonymous):

what proportion of a normal distribution falls between z=-.65 and z=+.65?

OpenStudy (anonymous):

look up the probability to the left of z=.65, and subtract the probability to the left of z=-.65.

OpenStudy (anonymous):

have you used a standard normal chart?

OpenStudy (anonymous):

yes but its confusing its my first time and i dont know where to start

OpenStudy (anonymous):

it is symmetric so you can also find 1-2* P(z<-.65)

OpenStudy (anonymous):

ok let me find a chart online and ill help you

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

http://www.stat.ucla.edu/~ywu/teaching/normal.pdf

OpenStudy (anonymous):

that's a good one, go there and let me know if the link works

OpenStudy (anonymous):

the link wont work for some reason

OpenStudy (anonymous):

ok let me see if I can find a similar one

OpenStudy (anonymous):

that's only a partial table but you can get the basic idea

OpenStudy (anonymous):

does it work?

OpenStudy (anonymous):

yes it works so do i look up -,65 and .65

OpenStudy (anonymous):

you would if it was a full table but we only have positive values here so we have to do some more work

OpenStudy (anonymous):

you can start by looking up .65

OpenStudy (anonymous):

the tenths place is in the first column, so go to .6

OpenStudy (anonymous):

ok so i see .6

OpenStudy (anonymous):

as you move over to the right you are adding one to the hundreths place. you can see the hundreths place an the top row of the chart

OpenStudy (anonymous):

0.7422

OpenStudy (anonymous):

look for .05, and find where the row .6 meets the column .05, that is the probability to the left of .65

OpenStudy (anonymous):

yep that's right.

OpenStudy (anonymous):

now we need to subtract P(z<-.65)

OpenStudy (anonymous):

standard normal is symmetric so P(z<-.65)=P(z>.65)

OpenStudy (anonymous):

so how would we find the negative

OpenStudy (anonymous):

P(z>.65)=1-P(z<.65)=1-.7422=.2578

OpenStudy (anonymous):

so P(z<.65)=.7422 P(z<-.65)=.2578 and we subtract them to find the probability in between .7422-.2578=.4844

OpenStudy (anonymous):

if you had a full table you could have just looked up z=-.65

OpenStudy (anonymous):

thanls so much, i really understand that now

OpenStudy (anonymous):

great glad it helped :)

OpenStudy (anonymous):

cya

OpenStudy (anonymous):

bye:)